# ExGRG: Explicitly-Generated Relation Graph for Self-Supervised Representation Learning

### v1

@@ Abstract.
Self-supervised Learning, Graph Representation Learning, Node Classification, Graph Neural Networks
1. Introduction
@@
To alleviate the shortcomings of data augmentation, our work [^1] Our code and pre-trained models will be publicly available. proposes integrating additional appropriate information derived from the input graph. This aids in determining which data samples are candidates to share similar representations. Rather than the prevailing rigid binary determination of whether two points should share a similar representation or not, we embrace a soft, probabilistic evaluation of the degree of similarity. Given the inadequacy and misguidance of graph augmentations, alternative cues are incorporated to guide the determination of analogous representations and assess the relative significance of each pairwise comparison. To achieve this, we propose explicitly generating a higher-order compositional relation graph that pinpoints crucial relation pairs, serving as candidates for the invariance loss. We propose a comprehensive strategy for constructing this compositional relation graph, leveraging insights from three key sources, each contributing to its distinct relation graph: (i) Neighborhood Similarity in the Representation Space, (ii) Higher-Order Graph Encodings, encompassing the source graph adjacency alongside diverse positional and structural encodings (PSEs), and (iii) A Deep Clustering module. We enrich the learning process, elevating the acquired representation and mitigating the deficiencies associated with relying exclusively on augmentations. While we motivate and emphasize the efficacy of our model based on VICReg (Bardes et al., 2021), our approach can be seamlessly integrated with other SSL methods simply by modifying their invariance terms through the exact process.
@@
Graph Neural Networks (GNNs) have the inherent ability to capture label-related patterns due to the message-passing paradigm and the intrinsic characteristics of graph datasets. Hence, we construct a k-Nearest Neighbour (kNN) graph on the representations. Nodes in close proximity within the kNN graph exhibit a heightened likelihood of sharing the same label class. Therefore, this kNN graph serves as our first source to construct the compositional relation graph. PSEs, alongside the adjacency information, constitute our second primary source. Inspired by (Lee et al., 2022), we employ the original input adjacency to construct a relation graph. Particularly in homophilic source graphs, adjacent nodes commonly share the same label class. Employing appropriate PSEs, transformers yield results comparable to message-passing networks. When a pair of points shares similar positional and structural properties, it indicates their overall similarity, offering auxiliary guidance for forming meaningful representations. Moreover, we incorporate a deep clustering algorithm rooted in the optimal transport (Asano et al., 2019) as the third source for guiding the invariance term. This module partitions the points into learnable clusters. Subsequently, we guide the model to create identical representations for points having similar soft clustering assignment distributions. This module is trained jointly in an end-to-end manner with the encoder, resembling an EM-style optimization approach in action, where both deep components evolve simultaneously, reflecting a dynamic interplay.
@@
Two theoretical justifications support our proposal: the Laplacian Eigenmap (LE) perspective and the EM viewpoint. Firstly, adopting a spectral manifold learning perspective, the original solely augmentation-based VICReg objective can be formulated as an LE optimization (Balestriero and LeCun, 2022). The implicit augmentation-based relation graph results in disconnected islands corresponding to a rank deficiency of its Laplacian. We mitigate this through explicitly generating the relation graph, facilitating the connection of these islands. Secondly, SSL methods address an underlying two-step EM characterized by two implicit sets of variables (Chen and He, 2021). We distinctly and explicitly outline the representatives of these two steps. The Expectation step (E-step) leverages the learned representation to dynamically generate a relation graph, identifying candidates to promote similar representations. The Maximization step (M-step) refines the encoder by incorporating information from the determined relation graph. These two modules synergize, with the relation graph generator pinpointing candidate pairs for the invariance term and the encoder ensuring the enforcement of identical representation pairs. Our introduced objective function facilitates and stabilizes the joint optimization of these modules, enabling simultaneous execution of the two steps in a single gradient update. Following extensive experimentation across diverse graph datasets, we showcase the superior quality of the learned representation, consistently outperforming previous methods.
2. Related Work
2.1. Self-supervised Representation Learning
@@
Contrastive SSL (Misra and Maaten, 2020; Bromley et al., 1994; Hjelm et al., 2019; Chen et al., 2020a; Hadsell et al., 2006; Ye et al., 2019; Wu et al., 2018; Chen et al., 2020c) constructs positive pairs employing data augmentation to push their representations closer while pulling apart the negative pairs utilizing InfoNCE (van den Oord et al., 2018). These pairs could be constructed on a mini-batch (Chen et al., 2020b) or employing memory banks (He et al., 2020). Clustering-based SSL (Bautista et al., 2016; Yang et al., 2016; Xie et al., 2016; Huang et al., 2019; Zhuang et al., 2019; Caron et al., 2019; Asano et al., 2019; Yan et al., 2020) employs a notion of clustering to form the representations by pseudo-labels (Caron et al., 2018) or enforcing identical clustering assignments of augmented pairs instead of directly enforcing identical features (Caron et al., 2020). Knowledge distillation-based SSL leverages student-teacher architectures (Hinton et al., 2015; Grill et al., 2020; Chen and He, 2021; Gidaris et al., 2021; Grill et al., 2020; Gidaris et al., 2020), employing an Exponential Moving Average encoder and a Stop-Gradient (SG) mechanism. Other works maximize the mutual information of representations (Ermolov et al., 2021; Zbontar et al., 2021). Notably, VICReg (Bardes et al., 2021) decorrelates embedding dimensions to prevent preserving redundant information, and VICRegL (Bardes et al., 2022) employs local and global features, enforcing the invariance term for specific image patches.
2.2. Graph Representation Learning
@@
GNNs (Kipf and Welling, 2016a; Veličković et al., 2017; Hamilton et al., 2017; Xu et al., 2018) accept node features and sufficient annotated data to form representations by recursively aggregating the neighbor’s information. Yet, to truly utilize the abundant unlabelled graph data, graph SSL pre-train GNNs to construct node- and graph-level representations without costly labels (García-Durán and Niepert, 2017; Kipf and Welling, 2016b; Bojchevski and Günnemann, 2017). Augmentation-based methods (You et al., 2020; Peng et al., 2020; Hassani and Khasahmadi, 2020; Zhu et al., 2021a, b, 2020b; Thakoor et al., 2021; Zhu et al., 2020a) encourage the representations to be invariant to a specifically designed graph transformation. A BERT-inspired method (Devlin et al., 2018; Hu et al., 2019) masks features of graphs with special structures. DGI (Veličković et al., 2018) aligns a local graph patch with the global one by maximizing the mutual information (Hjelm et al., 2019), followed by edge and node feature extensions (Peng et al., 2020; Jing et al., 2021) and tackling graph classification (Sun et al., 2019). Contrastive methods (Zhu et al., 2020b, 2021b; You et al., 2020; Hassani and Khasahmadi, 2020; Tian et al., 2020; Liu et al., 2023b) are generally inspired by SimCLR (Chen et al., 2020b), employing negative pairs resulting in the sampling bias issue (Bielak et al., 2021), meaning some negative samples may have similar semantics to the anchor while being pulled apart. PGCL (Lin et al., 2022) tackles the sampling bias, constructing the negative pairs from different clusters, while BGRL (Thakoor et al., 2021) adopts the non-contrastive BYOL (Grill et al., 2020).
@@
3. Method
Figure 1. The comprehensive architecture of ExGRG. We incorporate augmented views $\mathcal{G}^{(1)}$ and $\mathcal{G}^{(2)}$ from source graph $\mathcal{G}^{(s)}$ to compose an input graph $\mathcal{G}$, which is then fed into encoder $f_{\theta}$ and expander $g_{\phi}$ to produce representations ${\bm{H}}$ and embeddings ${\bm{Z}}$. We explicitly generate a compositional relation graph ${\bm{G}}$, aggregated from various intermediate relation graphs, to guide the invariance term $\mathcal{L}_{I^{\prime}}$ instead of relying solely on augmentations.
3.1. Preliminary
3.1.1. Graph SSL
@@
The essence of SSL lies in employing signals inherent in data points as a form of supervision. SSL pre-training aims to discover representations of input points by identifying $f_{\theta}$ and $g_{\phi}$. Subsequently, the pre-trained $f_{\theta}$ is utilized in downstream tasks, predicting ${\bm{Y}}^{s}$ from $\mathcal{G}^{s}$, through linear probing, nonlinear probing, or fine-tuning. In linear probing, the prevailing approach in graph SSL, a single linear layer is affixed to the frozen pre-trained $f_{\theta}$ while detaching $g_{\phi}$. This linear layer is trained for the downstream node classification task, utilizing the frozen representations as inputs. Though ExGRG centers around a node-level representation learning task, it can be seamlessly extended to learn graph- and link-level representations, with details provided in Appendix A.1.
3.1.2. VICReg and Relation Graph
Contrastive SSL prevents the representation collapse by incorporating negative pairs (Chen et al., 2020b). In contrast, the VICReg objective $\mathcal{L}_{Vic}$ (Eq. 3) introduces constraints to maximize the volume of the embedding space through two regularization terms, variance $\mathcal{L}_{V}$ and covariance $\mathcal{L}_{C}$ (Eq. 1). The term $\mathcal{L}_{V}$ ensures variance of embedding dimensions exceeds a threshold, while $\mathcal{L}_{C}$ enforces decorrelation between every pair of dimensions. The invariance term $\mathcal{L}_{I}$ (Eq. 2) encourages similar representations for semantically related points through augmentation.
(1) $\begin{split}\mathcal{L}_{V}=\sum_{k=1}^{D_{Z}}\max\left(0,1-\sqrt{\mathrm{Cov}({\bm{Z}})_{k,k}}\right),\\ \mathcal{L}_{C}=\sum_{k=1}^{D_{Z}}\sum_{j=1,j\not=k}^{D_{Z}}\mathrm{Cov}({\bm{Z}})^{2}_{k,j}.\end{split}$
(2) $\mathcal{L}_{I}=\sum_{(i,j)\in S_{a}}\|{\bm{Z}}_{i,.}-{\bm{Z}}_{j,.}\|_{2}^{2}$
(3) $\begin{split}\mathcal{L}_{Vic}=\alpha\mathcal{L}_{V}+\beta\mathcal{L}_{C}+\gamma\mathcal{L}_{I}\end{split}$
@@ 3.1.2. VICReg and Relation Graph
3.2. Explicit Compositional Relation Graph
@@
3.2.1. Laplacian Eigenmap Perspective
In variance-covariance-constrained VICReg, $\alpha=\beta\gg\gamma$, signifying that the variance and covariance terms in $\mathcal{L}_{Vic}$ are already satisfied (Eq. 3). Optimizing this VICReg objective is equivalent to solving the LE problem in Appendix Eq. 14 (Balestriero and LeCun, 2022). The distinction lies in VICReg’s ${\bm{G}}$ being determined through data augmentations, while in LE, it can be designed by forming kNN graphs in the input. Due to the duality between $S_{a}$ and $G^{a}$, the augmentation-based relation graph $G^{a}$ consists of $M=\frac{N}{2}$ disconnected islands. Therefore, the corresponding Laplacian $L^{a}$ for ${\bm{G}}^{a}$ suffers from rank deficiency, as it has $\frac{N}{2}$ zero eigenvalues. In addition, due to the shortcomings of graph augmentations, connections between pairs of data points within each island in ${\bm{G}}^{a}$ tend to be much less valuable compared to the vision domain. To address this challenge, we mitigate the rank deficiency of $L^{a}$ by constructing a more informative ${\bm{G}}$ through the incorporation of additional sources to guide the invariance term. We introduce meaningful entries to ${\bm{G}}$, establishing connections among the disconnected islands while assigning soft importance to each connection.
3.2.2. Alternating Two-Step EM Perspective
SimSiam (Chen and He, 2021) hypothesizes that SSL methods with Siamese architectures share an underlying optimization problem that could be effectively modeled employing the EM framework. In this optimization, SSL pre-training deals with two implicit sets of variables, reflecting an implied alternating optimization for each set in EM. Employing SG emerges as a viable approach to prevent collapse, resembling the alternating two-step EM optimization style in a single gradient update. The E-step is performed utilizing a predictor. Instead of directly comparing augmented views, the predictor assigns implicit targets for each anchor point, upon which the invariance loss is applied. In the M-step, the encoder is updated but only through one of the two available paths employing the SG technique.
@@ 3.2.2. Alternating Two-Step EM Perspective
3.2.3. Intra-view and Soft Relations
@@
3.3. Strategies for Constructing Relation Graphs
@@
3.3.1. Neighborhood Similarity in the Representation Space
@@
3.3.2. Higher-Order Graph Encodings
In homophilic input graphs, the proximity of two nodes in $\mathcal{G}^{s}$ indicates sharing the same class label in ${\bm{Y}}^{s}$ (Zhu et al., 2020c). Therefore, depending on the degree of homophily in the source graph $\mathcal{G}^{s}$, the edges between source points in ${\bm{A}}^{s}$ can be exploited to construct an informative relation matrix ${\bm{G}}^{A}=F_{A}({\bm{A}}^{s})\in\mathbb{R}^{N\times N}$, where $F_{A}$ extracts the adjacency information corresponding to the $N$ points in the mini-batch. Recognizing that ${\bm{G}}^{k}$ or ${\bm{G}}^{A}$ are often noisy on their own (Lee et al., 2022), we propose another more conservative filtered version through element-wise multiplication as ${\bm{G}}^{A^{\prime}}={\bm{G}}^{A}\odot{\bm{G}}^{k}$. This ${\bm{G}}^{A^{\prime}}$ retains the soft similarities of ${\bm{G}}^{k}$, with additional details provided in Appendix § A.5.

### v2

@@ Abstract
1 Introduction
@@
To alleviate the shortcomings of data augmentation, our work proposes integrating additional appropriate information derived from the input graph. This aids in determining which data samples are candidates to share similar representations. Rather than the prevailing rigid binary determination of whether two points should share a similar representation or not, we embrace a soft, probabilistic evaluation of the degree of similarity. Given the inadequacy and misguidance of graph augmentations, alternative cues are incorporated to guide the determination of analogous representations and assess the relative significance of each pairwise comparison. To achieve this, we propose explicitly generating a higher-order compositional relation graph that pinpoints crucial relation pairs, serving as candidates for the invariance loss. We propose a comprehensive strategy for constructing this compositional relation graph, leveraging insights from three key sources, each contributing to its distinct relation graph: (i) Neighborhood Similarity in the Representation Space, (ii) Higher-Order Graph Encodings, encompassing the source graph adjacency alongside diverse positional and structural encodings (PSEs), and (iii) A Deep Clustering module. We enrich the learning process, elevating the acquired representation and mitigating the deficiencies associated with relying exclusively on augmentations. While we motivate and emphasize the efficacy of our model based on VICReg (Bardes et al., 2021), our approach can be seamlessly integrated with other SSL methods simply by modifying their invariance terms through the exact process.
@@
Graph Neural Networks (GNNs) have the inherent ability to capture label-related patterns due to the message-passing paradigm and the intrinsic characteristics of graph datasets. Hence, we construct a k-Nearest Neighbour (kNN) graph on the representations. Nodes in close proximity within the kNN graph exhibit a heightened likelihood of sharing the same label class. Therefore, this kNN graph serves as our first source to construct the compositional relation graph. PSEs, alongside the adjacency information, constitute our second primary source. Inspired by Lee et al. (2022), we employ the original input adjacency to construct a relation graph. Particularly in homophilic source graphs, adjacent nodes commonly share the same label class. Employing appropriate PSEs, transformers yield results comparable to message-passing networks. When a pair of points shares similar positional and structural properties, it indicates their overall similarity, offering auxiliary guidance for forming meaningful representations. Moreover, we incorporate a deep clustering algorithm rooted in the optimal transport (Asano et al., 2019) as the third source for guiding the invariance term. This module partitions the points into learnable clusters. Subsequently, we guide the model to create identical representations for points having similar soft clustering assignment distributions. This module is trained jointly in an end-to-end manner with the encoder, resembling an EM-style optimization approach in action, where both deep components evolve simultaneously, reflecting a dynamic interplay.
@@
Two theoretical justifications support our proposal: the Laplacian Eigenmap (LE) perspective and the EM viewpoint. Firstly, adopting a spectral manifold learning perspective, the original solely augmentation-based VICReg objective can be formulated as an LE optimization (Balestriero & LeCun, 2022). The implicit augmentation-based relation graph results in disconnected islands corresponding to a rank deficiency of its Laplacian. We mitigate this through explicitly generating the relation graph, facilitating the connection of these islands. Secondly, SSL methods address an underlying two-step EM characterized by two implicit sets of variables (Chen & He, 2021). We distinctly and explicitly outline the representatives of these two steps. The Expectation step (E-step) leverages the learned representation to dynamically generate a relation graph, identifying candidates to promote similar representations. The Maximization step (M-step) refines the encoder by incorporating information from the determined relation graph. These two modules synergize, with the relation graph generator pinpointing candidate pairs for the invariance term and the encoder ensuring the enforcement of identical representation pairs. Our introduced objective function facilitates and stabilizes the joint optimization of these modules, enabling simultaneous execution of the two steps in a single gradient update. Following extensive experimentation across diverse graph datasets, we showcase the superior quality of the learned representation, consistently outperforming previous methods.
2 Related Work
2.1 Self-supervised Representation Learning
@@
Contrastive SSL (Misra & Maaten, 2020; Bromley et al., 1994; Hjelm et al., 2019; Chen et al., 2020c; Hadsell et al., 2006; Ye et al., 2019; Wu et al., 2018; Chen et al., 2020b) constructs positive pairs employing data augmentation to push their representations closer while pulling apart the negative pairs utilizing InfoNCE (van den Oord et al., 2018). These pairs could be constructed on a mini-batch (Chen et al., 2020a) or employing memory banks (He et al., 2020). Clustering-based SSL (Bautista et al., 2016; Yang et al., 2016; Xie et al., 2016; Huang et al., 2019; Zhuang et al., 2019; Caron et al., 2019; Asano et al., 2019; Yan et al., 2020) employs a notion of clustering to form the representations by pseudo-labels (Caron et al., 2018) or enforcing identical clustering assignments of augmented pairs instead of directly enforcing identical features (Caron et al., 2020). Knowledge distillation-based SSL leverages student-teacher architectures (Hinton et al., 2015; Grill et al., 2020; Chen & He, 2021; Gidaris et al., 2021; Grill et al., 2020; Gidaris et al., 2020), employing an Exponential Moving Average encoder and a Stop-Gradient (SG) mechanism. Other works maximize the mutual information of representations (Ermolov et al., 2021; Zbontar et al., 2021). Notably, VICReg (Bardes et al., 2021) decorrelates embedding dimensions to prevent preserving redundant information, and VICRegL (Bardes et al., 2022) employs local and global features, enforcing the invariance term for specific image patches.
2.2 Graph Representation Learning
@@
GNNs (Kipf & Welling, 2016a; Veličković et al., 2017; Hamilton et al., 2017; Xu et al., 2018) accept node features and sufficient annotated data to form representations by recursively aggregating the neighbor’s information. Yet, to truly utilize the abundant unlabelled graph data, graph SSL pre-train GNNs to construct node- and graph-level representations without costly labels (García-Durán & Niepert, 2017; Kipf & Welling, 2016b; Bojchevski & Günnemann, 2017). Augmentation-based methods (You et al., 2020; Peng et al., 2020; Hassani & Khasahmadi, 2020; Zhu et al., 2021a; b; 2020c; Thakoor et al., 2021; Zhu et al., 2020b) encourage the representations to be invariant to a specifically designed graph transformation. A BERT-inspired method (Devlin et al., 2018; Hu et al., 2019) masks features of graphs with special structures. DGI (Veličković et al., 2018) aligns a local graph patch with the global one by maximizing the mutual information (Hjelm et al., 2019), followed by edge and node feature extensions (Peng et al., 2020; Jing et al., 2021) and tackling graph classification (Sun et al., 2019). Contrastive methods (Zhu et al., 2020c; 2021b; You et al., 2020; Hassani & Khasahmadi, 2020; Tian et al., 2020; Liu et al., 2023b) are generally inspired by SimCLR (Chen et al., 2020a), employing negative pairs resulting in the sampling bias issue (Bielak et al., 2021), meaning some negative samples may have similar semantics to the anchor while being pulled apart. PGCL (Lin et al., 2022) tackles the sampling bias, constructing the negative pairs from different clusters, while BGRL (Thakoor et al., 2021) adopts the non-contrastive BYOL (Grill et al., 2020).
@@
3 Method
Figure 1: The comprehensive architecture of ExGRG. We incorporate augmented views $\mathcal{G}^{(1)}$ and $\mathcal{G}^{(2)}$ from source graph $\mathcal{G}^{(s)}$ to compose an input graph $\mathcal{G}$, which is then fed into encoder $f_{\theta}$ and expander $g_{\phi}$ to produce representations ${\bm{H}}$ and embeddings ${\bm{Z}}$. We explicitly generate a compositional relation graph ${\bm{G}}$, aggregated from various intermediate relation graphs, to guide the invariance term $\mathcal{L}_{I^{\prime}}$ instead of relying solely on augmentations.
3.1 Preliminary
3.1.1 Graph SSL
@@
The essence of SSL lies in employing signals inherent in data points as a form of supervision. SSL pre-training aims to discover representations of input points by identifying $f_{\theta}$ and $g_{\phi}$. Subsequently, the pre-trained $f_{\theta}$ is utilized in downstream tasks, predicting ${\bm{Y}}^{s}$ from $\mathcal{G}^{s}$, through linear probing, nonlinear probing, or fine-tuning. In linear probing, the prevailing approach in graph SSL, a single linear layer is affixed to the frozen pre-trained $f_{\theta}$ while detaching $g_{\phi}$. This linear layer is trained for the downstream node classification task, utilizing the frozen representations as inputs. Though ExGRG centers around a node-level representation learning task, it can be seamlessly extended to learn graph- and link-level representations, with details provided in Appendix § A.1.
3.1.2 VICReg and Relation Graph
Contrastive SSL prevents the representation collapse by incorporating negative pairs (Chen et al., 2020a). In contrast, the VICReg objective $\mathcal{L}_{Vic}$ (Eq. 3) introduces constraints to maximize the volume of the embedding space through two regularization terms, variance $\mathcal{L}_{V}$ and covariance $\mathcal{L}_{C}$ (Eq. 1). The term $\mathcal{L}_{V}$ ensures variance of embedding dimensions exceeds a threshold, while $\mathcal{L}_{C}$ enforces decorrelation between every pair of dimensions. The invariance term $\mathcal{L}_{I}$ (Eq. 2) encourages similar representations for semantically related points through augmentation.
$\begin{split}\mathcal{L}_{V}=\sum_{k=1}^{D_{Z}}\max\left(0,1-\sqrt{\mathrm{Cov}({\bm{Z}})_{k,k}}\right),\\ \mathcal{L}_{C}=\sum_{k=1}^{D_{Z}}\sum_{j=1,j\not=k}^{D_{Z}}\mathrm{Cov}({\bm{Z}})^{2}_{k,j}.\end{split}$ (1)
$\mathcal{L}_{I}=\sum_{(i,j)\in S_{a}}\|{\bm{Z}}_{i,.}-{\bm{Z}}_{j,.}\|_{2}^{2}$ (2)
$\begin{split}\mathcal{L}_{Vic}=\alpha\mathcal{L}_{V}+\beta\mathcal{L}_{C}+\gamma\mathcal{L}_{I}\end{split}$ (3)
@@ 3.1.2 VICReg and Relation Graph
3.2 Explicit Compositional Relation Graph
@@
3.2.1 Laplacian Eigenmap Perspective
In variance-covariance-constrained VICReg, $\alpha=\beta\gg\gamma$, signifying that the variance and covariance terms in $\mathcal{L}_{Vic}$ are already satisfied (Eq. 3). Optimizing this VICReg objective is equivalent to solving the LE problem in Appendix Eq. 14 (Balestriero & LeCun, 2022). The distinction lies in VICReg’s ${\bm{G}}$ being determined through data augmentations, while in LE, it can be designed by forming kNN graphs in the input. Due to the duality between $S_{a}$ and $G^{a}$, the augmentation-based relation graph $G^{a}$ consists of $M=\frac{N}{2}$ disconnected islands. Therefore, the corresponding Laplacian $L^{a}$ for ${\bm{G}}^{a}$ suffers from rank deficiency, as it has $\frac{N}{2}$ zero eigenvalues. In addition, due to the shortcomings of graph augmentations, connections between pairs of data points within each island in ${\bm{G}}^{a}$ tend to be much less valuable compared to the vision domain. To address this challenge, we mitigate the rank deficiency of $L^{a}$ by constructing a more informative ${\bm{G}}$ through the incorporation of additional sources to guide the invariance term. We introduce meaningful entries to ${\bm{G}}$, establishing connections among the disconnected islands while assigning soft importance to each connection.
3.2.2 Alternating Two-Step EM Perspective
SimSiam (Chen & He, 2021) hypothesizes that SSL methods with Siamese architectures share an underlying optimization problem that could be effectively modeled employing the EM framework. In this optimization, SSL pre-training deals with two implicit sets of variables, reflecting an implied alternating optimization for each set in EM. Employing SG emerges as a viable approach to prevent collapse, resembling the alternating two-step EM optimization style in a single gradient update. The E-step is performed utilizing a predictor. Instead of directly comparing augmented views, the predictor assigns implicit targets for each anchor point, upon which the invariance loss is applied. In the M-step, the encoder is updated but only through one of the two available paths employing the SG technique.
@@ 3.2.2 Alternating Two-Step EM Perspective
3.2.3 Intra-view and Soft Relations
@@
3.3 Strategies for Constructing Relation Graphs
@@
3.3.1 Neighborhood Similarity in the Representation Space
@@
3.3.2 Higher-Order Graph Encodings
In homophilic input graphs, the proximity of two nodes in $\mathcal{G}^{s}$ indicates sharing the same class label in ${\bm{Y}}^{s}$ (Zhu et al., 2020a). Therefore, depending on the degree of homophily in the source graph $\mathcal{G}^{s}$, the edges between source points in ${\bm{A}}^{s}$ can be exploited to construct an informative relation matrix ${\bm{G}}^{A}=F_{A}({\bm{A}}^{s})\in\mathbb{R}^{N\times N}$, where $F_{A}$ extracts the adjacency information corresponding to the $N$ points in the mini-batch. Recognizing that ${\bm{G}}^{k}$ or ${\bm{G}}^{A}$ are often noisy on their own (Lee et al., 2022), we propose another more conservative filtered version through element-wise multiplication as ${\bm{G}}^{A^{\prime}}={\bm{G}}^{A}\odot{\bm{G}}^{k}$. This ${\bm{G}}^{A^{\prime}}$ retains the soft similarities of ${\bm{G}}^{k}$, with additional details provided in Appendix § A.5.